A sample of carbon dioxide (CO2) gas has a volume of 15.2 L at a pressure of 1.35 atm and a



temperature of 33°C. Determine the following for this gas sample.



a. Volume, in liters, at T= 35°C and P= 3.50 atm



b. Pressure, in atmospheres, T= 42°C and V= 10.0 L



c. Temperature, in degrees Celsius, at P= 7.00 atm and V= 0.973 L



d. Volume, in milliliters, at T= 97°C and P= 6.70 atm

1
Expert's answer
2021-11-19T14:06:02-0500

1)using general gas law

(P1V1/T1)=(P2V2/T2)

V2=P1V1T2/P2T1

V2=(1.35×15.2×308)/3.5×306

V2=5.90L

2)P2=P1V1T2/V2T1

P2=(1.35×15.2×315)/10×306

P2=2.11atm

3)T2=P2V2T1/P1V1

T2=(7×0.973×306)/1.3×15.2

T2=105.47K

4)V2=P1V1T2/P2T1

V2=(1.35×15.2×370)/6.7×306

V2=3.7L

V2=0.0037mL